Re-write the quadratic function below in Standard Form
y = 9(x + 2)^2 + 8

Answer by ikleyn(52846)

.

y = 9(x + 2)^2 + 8 = 9*(x^2 + 4x + 4) + 8 = 9x^2 + 36x + 36 + 8 = 9x^2 + 36x + 44.


That's all. Very easy exercise.